### v2

@@ Abstract
As the volume of scientific literature continues to grow, efficient knowledge organization is an increasingly challenging task. Traditional structuring of scientific content is time-consuming and requires significant domain expertise, increasing the need for tool support. Our goal is to create a Human-in-the-Loop (HITL) workflow that supports researchers in creating and structuring scientific knowledge, leveraging neural models and knowledge graphs, exemplified using the Open Research Knowledge Graph (ORKG). The workflow aims to automate key steps, including data extraction and knowledge structuring, while keeping user oversight through human validation. We developed a modular framework implementing the workflow and evaluated it along the Quality Improvement Paradigm (QIP) with participants from the ORKG user community. The evaluation indicated that the framework is highly usable and provides practical support. It significantly reduces the time and effort required to transition from a research interest to literature-based answers by streamlining the import of information into a knowledge graph. Participants evaluated the framework with an average System Usability Scale (SUS) score of 84.17, an A+ – the highest achievable rating. They also reported that it improved their time spent, previously between 4 hours and two weeks, down to an average of 24:40 minutes. The tool streamlines the creation of scientific corpora and extraction of structured knowledge for KG integration by leveraging LLMs and user-defined models, significantly accelerating the review process. However, human validation remains essential throughout the extraction process, and future work is needed to improve extraction accuracy and entity linking to existing knowledge resources.
@@
HITL Knowledge Graph Literature Review LLM Modular framework Knowledge Organization
@@ 1 Introduction
Conducting literature reviews is indispensable for scientific research. This task is inherently challenging, as it requires a significant amount of time and effort to gain an overview of the existing research and knowledge on a specific topic, helping to support methodology, contextualize research, or identify gaps in knowledge [33]. With the rapid growth in scientific discovery and distribution of knowledge, it has become even more difficult to produce high-quality reviews [23]. Recent advancements in artificial intelligence (AI), particularly the rise of large language models (LLMs) and AI-based Literature Reviews (AILR) [39], offer promising opportunities to streamline and automate repeatable and time-consuming processes [29]. However, despite their growing popularity, AI models face significant limitations, primarily the tendency to produce hallucinations, resulting in inconsistencies and confusion [28]. Combining AI capabilities with a consistent knowledge base, called neuro-symbolic AI, mitigates these risks by leveraging reliable knowledge representations to alleviate AI’s unreliability [3]. The knowledge bases need to be Findable, Accessible, Interoperable, and Reusable (FAIR) [40] to address the requirements of the scholarly domain and AI technology [2]. Such solutions already exist, like the Open Research Knowledge Graph (ORKG) [^1] https://orkg.org/ and its neuro-symbolic service Ask [3], offering a novel approach by representing scientific information in a structured format, providing both human- and machine-readability. However, organizing knowledge at scale reintroduces the original challenge of requiring human oversight, leading to manual work and a circular dependency. To address this issue, an iterative, Human-in-the-Loop (HITL), AI-assisted knowledge organization approach is needed, constantly improving scholarly knowledge representation.
Our contributions include 1) A HITL workflow, combining LLM support with Natural Language Processing (NLP) techniques to assist users by streamlining, automating, and formalizing tasks of literature reviews. 2) A modular framework implementing this workflow to build a scientific corpus of literature and transform information into structured knowledge, ready for integration into knowledge graphs. 3) An iterative evaluation with nine researchers over 4 months, developing and validating the frameworks’ capacity to ensure greater efficiency while allowing users to maintain oversight of the automated processes.
This work is structured as follows: In section 2 we give an overview of the background and related work in the HITL context. Our approach is described in section 3 , followed by an outline of the implementation in section 4. We evaluate the developed solution in section 5 and discuss the results in section 6. We conclude our work in section 7.
@@ 2 Background and Related Work
Knowledge Organization. Knowledge Graphs (KGs) are designed to represent and interconnect information in a structured, machine-readable format [36], like the scientific knowledge graph (SKG) Semantic Scholar [38]. Manually preparing data for KGs, particularly SKGs, is time-consuming and lacks scalability, as it requires domain specialists to search, interpret, and structure scientific literature [12]. Without domain experts, there is a higher risk of integrating incorrect, incomplete, or outdated information into the SKG, reducing its reliability and quality [15]. The quality of SKGs heavily depends on their construction method, with significantly different results between manual and automated creation methods [13]. The ORKG is a versatile scientific knowledge platform that uses crowdsourcing to organize and link scientific knowledge across various topics and domains [3]. Karras et al. [18, 19, 20] demonstrated how the ORKG can enhance and ensure FAIR [40] literature reviews, highlighting their importance in providing research with structured, long-term available data. The ORKG Assistant for Scientific Knowledge (Ask) [3] applies a neuro-symbolic approach using semantic search, LLMs, and KGs. It exemplifies a next-generation scholarly search and exploration system, combining semantic knowledge representation with AI.
Artificial Intelligence (AI). AI and NLP techniques can automate content generation and assignment [10], but these methods often lack transparency and accountability while struggling with semantic context [21]. Fully relying on AI-supported technologies poses significant risks, making human supervision and validation indispensable [30]. The study conducted by Danler et al. [9] demonstrates that, while AI-generated outputs can be of high quality, “relying solely on AI-generated outputs for crafting a scientific paper is not advisable” [9, p. 207]. Bacinger et al. [5] present Litre Assistant, a machine-learning-based solution to determining a corpus of scientific papers as the initial task of the Systematic Literature Review (SLR) process. Their work introduces a system that allows users to iteratively select desired papers as a set of training data for machine learning (ML) algorithms for an automated classification process. Ye et al. [43] propose a hybrid, “human-oriented”, but AI-based workflow to support specific steps of an SLR. Their workflow screens full-text articles to extract relevant data and allows the user to perform inclusion and exclusion decisions, resulting in a spreadsheet of LLM-generated data. The work of Alshami et al. [1] introduces a method where humans review ChatGPT-generated data for literature setup, abstract screening, data extraction, and analysis. However, relying solely on ChatGPT restricts access to real-time data, limiting the research scope to the model’s training data and reducing its applicability.
@@
Our work builds on the discussed approaches, emphasizing the role of HITL in the context of AI-assisted knowledge graph construction. In this spirit, we aim to develop a solution that supports transforming completely unstructured input into knowledge graph representations, with a focus on efficiency and usability. These knowledge graphs will then provide the human- and machine-readable context required by neuro-symbolic AI to sustainably address the challenge of scholarly knowledge organization.
@@ 3 Approach
We adopt the Quality Improvement Paradigm (QIP) [7] to guide the iterative development of a modular framework, implementing a HITL workflow, for transforming scientific corpora into structured knowledge. The approach was developed in close collaboration with ORKG curators, researchers from diverse domains funded through grants to create structured comparisons, a key content type presenting tabular overviews of state-of-the-art research in the ORKG [26]. While their financial affiliation may introduce some bias, their interest in improving tools that support their contributions fosters practical, needs-driven feedback. An overview of the development process is shown in Figure 1, with detailed tasks in LABEL:tab:qip_stepsLABEL:tab:qip_steps. The Project Learning cycle focused on short-term usability improvements based on curator feedback, while the Corporate Learning cycle addressed long-term enhancements to ORKG infrastructure beyond the project.
Figure 1: Application of the Quality Improvement Paradigm (QIP) methodology with ORKG curators. Derived from [27].
Table 1: QIP step descriptions
@@
3. Choose process We selected external ORKG curators as a focus group due to their research diversity and active involvement in literature-based knowledge graph curation. Using an iterative participatory design approach with bi-weekly meetings, we co-developed modular Jupyter Notebooks tailored to flexible workflow stages (cf. subsection 3.2).
@@
6. Package Finally, we compiled reusable artifacts, including framework components, evaluation metrics, and key insights, to support broader adoption. We also assessed ORKG’s infrastructure to enable sustainable integration (cf. sections 5.0.1 and 6).
@@ 3.1 Requirements
To define functional and non-functional requirements (see LABEL:tab:requirementsLABEL:tab:requirements) for our HITL framework, we drew inspiration from the work of Wittenborg et al. [41], who specified requirements for the typical phases of an SLR. While our framework does not aim to fully replicate or enforce the rigor of a formal SLR, we adopt this process-oriented perspective to ensure relevance and alignment with real-world research workflows. In collaboration with the curators and domain experts behind SWARM-SLR, we adapted these ideas to guide the design of our system requirements and to support a more lightweight, customizable process for literature exploration. Our goal is to offer a modular framework that enables the generation and processing of a literature corpus without imposing the complexity of a full SLR pipeline. To ensure modularity and usability, we divide the workflow into two main stages, each addressing a distinct set of user needs. The framework is designed to allow users to begin directly with the second stage if a relevant corpus has already been assembled.
@@
FR7: The framework must allow a user to use an existing corpus to proceed with knowledge extraction. Partial
@@ 3.2 Workflow
In the following, we detail how our workflow is designed to address the defined requirements outlined above. While not aiming to fully replicate the SLR process, the two workflow stages are broadly inspired by its typical phases and serve as a practical structure for supporting key tasks. Our proposed HITL workflow blends automation with human validation, allowing users to apply their expertise where it’s most valuable, especially in interpreting and refining scientific content. To reduce the manual burden, we focus on automating the most time-consuming steps, while still giving users control over key decisions. Because the workflow can begin with any research interest, no pre-structured input is needed, making the process flexible and adaptable to a wide range of review scenarios.
Stage 1: Creating a scientific corpus of literature.
The first stage of the workflow is dedicated to creating a corpus of scientific publications relevant to the user’s research interest, as depicted in Figure 2. Users begin by defining a research interest, which can be used by the LLM to suggest relevant keywords, though this step is optional. The user then manually formulates a search query, which is used to search selected sources. Search parameters, such as output size and the inclusion of only openly accessible full texts, can be further configured, as a subset of users prefers a broader range of materials and does not prioritize full-text access. The retrieved papers are then reviewed and selectively chosen by the user, forming the input for Stage 2.
@@
Stage 2: Knowledge extraction and validation.
The second stage processes the knowledge within the corpus, as shown in Figure 3. The user begins by defining the properties of the data model to be extracted from each paper. The LLM agent then parses and interprets the corpus, generating output based on the specified properties. If any information is not found, the LLM flags this to mitigate the risk of hallucination. The user can manually refine, revise, and modify the results to ensure accuracy, ultimately validating the output. The final results are formatted, and optionally, entity extraction is performed to create output that is suitable for KG integration.
@@ 4 Implementation
We implement the described functions as a modular framework in two Jupyter Notebooks using Python (NR5). These notebooks include automated modules, visualizations, and browser-based interactive control elements (NR2). We leverage the interactive HTML widgets from the Python package ipywidgets as well as the IPython toolkit for abstracting complex code interactions and enhancing the usability and interactivity. Users can interact with UI elements like input fields and checkboxes for easy configuration. They can also extend functionality by adding query sources or modifying the UI, cells, and behavior as needed (NR1).
The first stage is designed to retrieve relevant papers based on a user-defined search query (FR2). Prior to this, the framework can suggest relevant keywords based on the user’s research interest (FR1) to support the formulation of an effective query. We utilize multiple APIs, including SemanticScholar, ArXiv and the SerpApi wrapper. Accessing diverse scholarly databases enables real-time search across various databases (FR3) based on user search preferences (FR4). These platforms return JSON objects with different data schemas, which we unify to ensure consistency (NR3). Most relevant elements, such as title, abstract, authors, publication date, and links to full texts are then displayed within the notebook in expandable accordions (FR5), as shown in Figure 4.
@@
Based on these suggestions, the user iteratively creates the corpus by selecting and downloading relevant papers (FR6). For reproducibility, the corresponding DOIs can be stored in a CSV file.
The second stage involves processing the corpus with the assistance of an LLM. While tools such as Ollama offer local access to multiple LLMs, we opted for a cloud-based solution that avoids reliance on local resources. We selected the Mistral-large-2 model, accessible via Mistral’s API and currently free of charge. This choice also enables specifying output formats as JSON, improving structure and compatibility with downstream processes. The tool prompts the LLM iteratively with separate instructions, including the defined properties by the user (FR8), and the pre-processed text extracted from each PDF in the corpus. The pre-processing pipeline begins by extracting raw text from each PDF using pdfminer. To avoid including external or irrelevant material in the prompt, the references section and appendices of each paper are identified and removed. Additionally, we apply regex-based heuristics to format and reconstruct the sentences and paragraphs, ensuring proper context is preserved for further processing. The generated output (FR9) is displayed to the user using a datagrid (FR11) shown in Figure 5. This allows users to review, edit, and decide which entries to include in the final CSV file. They can also refine the data model and initiate a new extraction process. However, previously generated extractions may be altered or lost, as the model reprocesses the data.
@@
We have begun to integrate entity linking (FR12) into our workflow as a step toward semantic enrichment. Following validation, the LLM-generated text is processed for entity extraction using Falcon 2.0 [31], a rule-based system capable of identifying and linking entities to knowledge graphs such as DBpedia [4] and Wikidata [37]. At this stage, we use Falcon 2.0 primarily for entity extraction. Further steps, such as full entity linking and seamless integration into a KG, are planned for future work to support more machine-readable enrichment of the extracted content. Previous work [17] addressed semantification of tables by aligning entities and structuring content during KG import.

### v3

@@ Abstract
As the volume of scientific literature grows, efficient knowledge organization becomes increasingly challenging. Traditional approaches to structuring scientific content are time-consuming and require significant domain expertise, highlighting the need for tool support. We present ExtracTable, a Human-in-the-Loop (HITL) workflow and framework that assists researchers in transforming unstructured publications into structured representations. The workflow combines large language models (LLMs) with user-defined schemas and is designed for downstream integration into knowledge graphs (KGs). Developed and evaluated in the context of the Open Research Knowledge Graph (ORKG), ExtracTable automates key steps such as document preprocessing and data extraction while ensuring user oversight through validation. In an evaluation with ORKG community participants following the Quality Improvement Paradigm (QIP), ExtracTable demonstrated high usability and practical value. Participants gave it an average System Usability Scale (SUS) score of 84.17 (A+, the highest rating). The time to progress from a research interest to literature-based insights was reduced from between 4 hours and 2 weeks to an average of 24:40 minutes. By streamlining corpus creation and structured data extraction for knowledge graph integration, ExtracTable leverages LLMs and user models to accelerate literature reviews. However, human validation remains essential to ensure quality, and future work will address improving extraction accuracy and entity linking to existing knowledge resources.
@@
HITL Knowledge Graph Integration Literature Review LLM Modular framework Knowledge Organization
@@ 1 Introduction
Conducting literature reviews is indispensable for scientific research. This task is inherently challenging, as it requires a significant amount of time and effort to gain an overview of the existing research and knowledge on a specific topic, helping to support methodology, contextualize research, or identify gaps in knowledge [34]. With the rapid growth in scientific discovery and distribution of knowledge, it has become even more difficult to produce high-quality reviews [24]. Recent advancements in artificial intelligence (AI), particularly the rise of large language models (LLMs) and AI-based Literature Reviews [39], offer promising opportunities to streamline and automate repeatable and time-consuming processes [30]. However, despite their growing popularity, AI models face significant limitations, primarily the tendency to produce hallucinations, resulting in inconsistencies and confusion [29]. Combining AI capabilities with a consistent knowledge base, called neuro-symbolic AI, mitigates these risks by leveraging reliable knowledge representations to alleviate AI’s unreliability [3]. The knowledge bases need to be Findable, Accessible, Interoperable, and Reusable (FAIR) [40] to address the requirements of the scholarly domain and AI technology [2]. Such solutions already exist, like the Open Research Knowledge Graph (ORKG) [^1] https://orkg.org/ and its neuro-symbolic service Ask [3], offering a novel approach by representing scientific information in a structured format, providing both human- and machine-readability. However, organizing knowledge at scale reintroduces the original challenge of requiring human oversight, leading to manual work and a circular dependency. To address this issue, an iterative, Human-in-the-Loop (HITL), AI-assisted knowledge organization approach is needed, constantly improving scholarly knowledge representation. Our contributions include 1) ExtracTable, a Human-in-the-Loop (HITL) approach comprising a workflow and modular implementation that combines large language model (LLM) support with Natural Language Processing (NLP) techniques to streamline, automate, and formalize literature review tasks. 2) The modular implementation of ExtracTable as a Jupyter-based framework to build scientific corpora and transform unstructured information into structured knowledge ready for knowledge graph integration. 3) An iterative evaluation with nine researchers over four months, demonstrating ExtracTable’s ability to improve efficiency while maintaining user oversight of automated processes. This work is structured as follows: In section 2 we give an overview of the background and related work in the HITL context. Our approach is described in section 3, followed by an outline of the implementation in section 4. We evaluate the developed solution of ExtracTable in section 5 and discuss the results in section 6. We conclude our work in section 7.
@@ 2 Background and Related Work
Knowledge Organization. Knowledge Graphs (KGs) are designed to represent and interconnect information in a structured, machine-readable format [36], like the scientific knowledge graph (SKG) Semantic Scholar [38]. Manually preparing data for KGs, particularly SKGs, is time-consuming and lacks scalability, as it requires domain experts to search, interpret, and structure scientific literature [12]. Without domain experts, there is a higher risk of integrating incorrect, incomplete, or outdated information into the SKG, reducing its reliability and quality [15]. The quality of SKGs heavily depends on their construction method, with significantly different results between manual and automated methods [13]. The ORKG is a versatile scientific knowledge platform that uses crowdsourcing to organize and link scientific knowledge across various topics and domains [3]. Karras et al. [21, 18, 19, 20] demonstrated how the ORKG can enhance and ensure FAIR [40] literature reviews, highlighting their importance in providing research with structured, long-term available data. The ORKG Assistant for Scientific Knowledge (Ask) [3] applies a neuro-symbolic approach using semantic search, LLMs, and KGs. It exemplifies a next-generation scholarly search and exploration system, combining semantic knowledge representation with AI.
Artificial Intelligence (AI). AI and NLP techniques can automate content generation and assignment [10], but these methods often lack transparency and accountability while struggling with semantic context [22]. Fully relying on AI-supported technologies poses significant risks, making human supervision and validation indispensable [31]. The study conducted by Danler et al. [9] demonstrates that, while AI-generated outputs can be of high quality, “relying solely on AI-generated outputs for crafting a scientific paper is not advisable” [9, p. 207]. Bacinger et al. [5] present Litre Assistant, a machine-learning-based solution to determining a corpus of scientific papers as the initial task of the Systematic Literature Review (SLR) process. Their work introduces a system that allows users to iteratively select desired papers as a set of training data for machine learning algorithms for an automated classification process. Ye et al. [43] propose a hybrid, “human-oriented”, but AI-based workflow to support specific steps of an SLR. Their workflow screens full-text articles to extract relevant data and allows the user to perform inclusion and exclusion decisions, resulting in a spreadsheet of LLM-generated data. The work of Alshami et al. [1] introduces a method where humans review ChatGPT-generated data for literature setup, abstract screening, data extraction, and analysis. However, relying solely on ChatGPT restricts access to real-time data, limiting the research scope to the model’s training data and reducing its applicability.
@@
Our work builds on the discussed approaches, emphasizing the role of HITL in the context of AI-assisted knowledge graph construction. In this spirit, we aim to develop a solution that transforms unstructured input into structured, human-validated outputs designed for downstream knowledge graph integration. Once integrated into knowledge graphs, these structured outputs can support the human- and machine-readable context needed by neuro-symbolic AI to sustainably advance scholarly knowledge organization.
@@ 3 Approach
We adopt the Quality Improvement Paradigm (QIP) [7] to guide the iterative development of ExtracTable, a Human-in-the-Loop (HITL) approach that includes a workflow and is realized as a modular framework for transforming scientific corpora into structured knowledge.
ExtracTable was developed in close collaboration with ORKG curators, researchers from diverse domains funded through grants to create structured comparisons, a key content type presenting tabular overviews of state-of-the-art research in the ORKG [27]. While their financial affiliation may introduce some bias, their interest in improving tools that support their contributions fosters practical, needs-driven feedback. An overview of the development process is shown in Figure 1, with detailed tasks in LABEL:tab:qip_stepsLABEL:tab:qip_steps. The Project Learning cycle focused on short-term usability improvements based on curator feedback, while the Corporate Learning cycle addressed long-term enhancements to ORKG infrastructure beyond the project.
Figure 1: Application of the Quality Improvement Paradigm (QIP) methodology to develop ExtracTable, together with ORKG curators. Derived from [28].
Table 1: QIP step descriptions for the development of ExtracTable
@@
3. Choose process We selected external ORKG curators as a focus group due to their research diversity and active involvement in literature-based knowledge graph curation. Using an iterative participatory design approach with bi-weekly meetings, we co-developed modular Jupyter Notebooks for ExtracTable tailored to flexible workflow stages (cf. subsection 3.2).
@@
6. Package Finally, we compiled reusable artifacts for ExtracTable, including framework components, evaluation metrics, and key insights, to support broader adoption. We also assessed ORKG’s infrastructure to enable sustainable integration (cf. sections 5.0.1 and 6).
@@ 3.1 Requirements
To define functional and non-functional requirements for ExtracTable (see LABEL:tab:requirementsLABEL:tab:requirements), we drew inspiration from Wittenborg et al. [41], who specified requirements for the typical phases of an SLR. While our framework does not aim to replicate or enforce the rigor of a formal SLR, we adopt this process-oriented perspective to ensure relevance and alignment with real-world research workflows. In collaboration with the SWARM-SLR curators and domain experts, we adapted these ideas to guide the design of system requirements and support a more lightweight, customizable process for literature exploration. Our goal is to offer a modular framework for ExtracTable that enables generation and processing of a literature corpus without the complexity of a full SLR pipeline. To ensure modularity and usability, the workflow is divided into two stages, and users can start with the second if a relevant corpus already exists.
@@
FR7: The framework must allow a user to use an existing corpus to proceed with knowledge extraction. Yes
@@ 3.2 Workflow
We now describe how the ExtracTable approach, centered on its HITL workflow, meets the requirements outlined above. While not aiming to replicate the full SLR process, the two workflow stages draw inspiration from its typical phases and provide a practical structure for key tasks. Our workflow blends automation with human validation, enabling users to focus their expertise on interpreting and refining scientific content. To reduce manual effort, we automate time-consuming steps while preserving user control over key decisions. The workflow can start from any research interest without requiring pre-structured input, making it flexible and adaptable to diverse review scenarios.
Stage 1: Creating a scientific corpus of literature. This stage focuses on building a corpus of publications relevant to the user’s research interest (see Figure 2). After defining a research interest and optionally using the LLM to suggest keywords, users formulate a search query to retrieve publications from selected sources. Search parameters, such as output size or restricting to openly accessible full texts, are configurable to suit user preferences. Retrieved papers are reviewed and selected by the user, forming the input for Stage 2.
@@
Stage 2: Knowledge extraction and validation. The second stage processes the knowledge within the corpus, as shown in Figure 3. The user begins by defining the properties of the data model to be extracted from each paper. The LLM agent then parses and interprets the corpus, generating output based on the specified properties. If any information is not found, the LLM flags this to mitigate the risk of hallucination. The user can manually refine, revise, and modify the results to ensure accuracy, ultimately validating the output. The final results are formatted, and optionally, entity linking is performed to create output that is suitable for KG integration as a machine-readable file.
@@ 4 Implementation
We implement the described functions as a modular framework for ExtracTable in two Jupyter Notebooks using Python (NR5). These notebooks provide automated modules, visualizations, and browser-rendered interactive controls (NR2). To enhance usability, we leverage interactive HTML widgets from ipywidgets and the IPython toolkit for abstracting complex code interactions. Users can configure workflows via UI elements such as input fields and checkboxes, and extend functionality by adding query sources or modifying the UI and behavior as needed (NR1). The implementation is modular and flexible, designed to support integration with various knowledge graphs.
The first stage retrieves relevant papers based on a user-defined search query (FR2). To support query formulation, the framework can optionally suggest keywords from the user’s research interest (FR1). We integrate multiple APIs, including SemanticScholar, ArXiv, and the SerpApi wrapper, enabling real-time searches across diverse scholarly databases (FR3) based on user preferences (FR4). Returned JSON objects with varying schemas are unified for consistency (NR3). Key elements such as title, abstract, authors, publication date, and full-text links are presented within expandable accordions in the notebook (FR5), as shown in Figure 4. Alternatively, users may work with an existing corpus, which can currently be imported from a Zotero collection (FR7).
@@
Based on these suggestions, the user iteratively creates the corpus by selecting and downloading relevant papers (FR6). For reproducibility, the corresponding DOIs can be exported to a CSV file.
The second stage processes the corpus using an LLM. We use the Mistral-large-2 model, accessible via Mistral’s API and currently free of charge. This model supports JSON-formatted outputs, enhancing structure and compatibility with downstream workflows. The tool prompts the LLM iteratively using user-defined properties (FR8) and pre-processed text extracted from each PDF. Pre-processing starts with raw text extraction via pdfminer. References and appendices are removed, and regex heuristics clean and reconstruct sentences and paragraphs to preserve context. Generated outputs (FR9) are displayed in a datagrid (FR11; see Figure 5), allowing users to review, edit, and select entries for final CSV export. Users can also refine the data model and rerun extraction, though previous results may be overwritten. An example of prompt and output can be found in the repository.
@@
We have begun integrating entity linking (FR12) as a step toward semantic enrichment. At this stage, we use Falcon 2.0[32] for entity extraction, identifying relevant concepts in free-text fields. Although Falcon already supports linking entities to knowledge graphs such as DBpedia[4] and Wikidata [37], we currently focus on extraction as a first step toward machine-actionability. This approach prepares the ground for entity linking and, ultimately, seamless integration with ORKG and other KGs in the future. Previous work [17] examined how knowledge representation evolves across formats and focused on semantically enriching tables by aligning entities and structuring content during KG import.